OpenStudy (anonymous):

Let the function f be defined by the equation y = f(x), where x and f(x) are real numbers. Find the domain of the function.

OpenStudy (anonymous):

\[f(x)=x+3 \over x-13\]

OpenStudy (anonymous):

all real x, \[x \neq13\]

OpenStudy (anonymous):

these are my choices for the answer a.(-[\infty\], 13) ∪ (13,[\infty\] ) b. (-[\infty\], -13) ∪ (13,[\infty\] ) c. (-[\infty\],[\infty\] ) d. (-[\infty\], 0] ∪ (13,[\infty\] ) e. (- [\infty\] , -13] ∪ [13,[\infty\] )

OpenStudy (anonymous):

suppose to be [\infty\] infinity sign

OpenStudy (anonymous):

then it would be "a" that includes all points except x=13

OpenStudy (anonymous):

thanks that where i thought you where going with it but was not sure..